# Spectral Graph Pruning Against Over-Squashing and Over-Smoothing

### v1

@@ Abstract
Message Passing Graph Neural Networks are known to suffer from two problems that are sometimes believed to be diametrically opposed: over-squashing and over-smoothing. The former results from topological bottlenecks that hamper the information flow from distant nodes and are mitigated by spectral gap maximization, primarily, by means of edge additions. However, such additions often promote over-smoothing that renders nodes of different classes less distinguishable. Inspired by the Braess phenomenon, we argue that deleting edges can address over-squashing and over-smoothing simultaneously. This insight explains how edge deletions can improve generalization, thus connecting spectral gap optimization to a seemingly disconnected objective of reducing computational resources by pruning graphs for lottery tickets. To this end, we propose a more effective spectral gap optimization framework to add or delete edges and demonstrate its effectiveness on large heterophilic datasets.
@@ 1 Introduction
Graphs are ubiquitous data structures that can model data from diverse fields ranging from chemistry (Reiser et al., 2022), biology (Bongini et al., 2023) to even high-energy physics (Shlomi et al., 2021). This has led to the development of deep learning techniques for graphs, commonly referred to as Graph Neural Networks (GNNs). The most popular GNNs follow the message-passing paradigm (Gori et al., 2005; Scarselli et al., 2009; Gilmer et al., 2017; Bronstein et al., 2021), where arbitrary differentiable functions, parameterized by neural networks, are used to diffuse information on the graph, consequently learning a graph-level representation. This representation can then be used for various downstream tasks like node classification, link prediction, and graph classification. Different types of GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Xu et al., 2019; Bodnar et al., 2021a, b; Bevilacqua et al., 2022), all tackling a variety of problems in various domains, have been proposed with varied degree of success. Despite their widespread use, GNNs have a number of inherent problems. These include limited expressivity, (Leman, 1968; Morris et al., 2019), over-smoothing (Li et al., 2019; NT & Maehara, 2019; Oono & Suzuki, 2020; Zhou et al., 2021), and over-squashing (Alon & Yahav, 2021; Topping et al., 2022).
@@
The phenomenon of over-squashing, first studied heuristically by Alon & Yahav (2021) and later theoretically formalized by Topping et al. (2022), is caused by the presence of structural bottlenecks in the graph. These bottlenecks can be attributed to the first non-zero eigenvalue of the normalized graph Laplacian, also known as the spectral gap. The smaller the gap, the more susceptible a graph is to over-squashing. Recent work has explored rewiring the input graph to address these bottlenecks. Some (Topping et al., 2022) focus on rewiring the input graph by calculating the discrete Ricci curvature, which is typically inefficient. Karhadkar et al. (2023) propose a computationally efficient first-order approximation of the spectral gap and use it as a criterion to add edges that maximize this quantity. While this is an effective strategy to mitigate over-squashing, it also exacerbates the problem of over-smoothing, where node features tend to converge to a non-informative limit because of repeated rounds of aggregation (Keriven, 2022). As a resolution that caters to both problems, we propose to leverage the Braess paradox (Braess, 1968; Eldan et al., 2017) that highlights certain edge deletions can also maximize the spectral gap. We propose to approximate the spectral change in a computationally efficient manner by leveraging Matrix Perturbation Theory (Stewart & Sun, 1990). Our proposed framework allows us to jointly address the problem of over-squashing, by increasing the spectral gap, and over-smoothing, by slowing down the rate of smoothing. We find that our method is especially effective in heterophilic graph settings, where we delete edges between nodes of different labels, thus preventing unnecessary aggregation. We empirically show that our proposed method outperforms other graph rewiring methods on node classification and graph classification tasks. We also show that spectral gap based edge deletions can help identify graph lottery tickets (GLTs) (Frankle & Carbin, 2019), that is, sparse sub-networks that can match the performance of dense networks. Graph pruning can destroy critical paths of information flow that enable effective message passing if it does not take the spectral gap of the resulting graph into account. Spectral gap based graph sparsification has the advantage that it does not depend on feature data or the GNN learning dynamics and could thus be performed at initialization, providing computational benefits from the start of training. We demonstrate that our proposal frequently outperforms state-of-the-art iterative magnitude pruning (Chen et al., 2021) that takes more information into account.
@@ 1.1 Contributions
1) Inspired by the Braess phenomenon, we gain theoretical insights into the potential of edge deletions to simultaneously reduce over-smoothing and over-squashing. 2) Leveraging matrix perturbation theory, we propose a Greedy graph pruning algorithm (ProxyDelete) that maximizes the spectral gap in a computationally efficient way. Similarly, our algorithm can also be utilized to add edges in a joint framework. We compare this approach with a novel graph rewiring scheme based on Eldan’s criterion (Eldan et al., 2017) that provides guarantees for edge deletions and a stopping criterion for pruning, but is computationally less efficient. 3) Our proposed graph modification strategy is capable of simultaneously addressing the problems of over-squashing and over-smoothing, especially in heterophilic settings. We demonstrate the effectiveness of our proposed framework on several semi-supervised node classification and graph classification tasks and show consistent improvements over existing baselines. 4) Our results connect literature on three seemingly disconnected topics: over-smoothing, over-squashing, and graph lottery tickets, which explain observed improvements in generalization performance by graph pruning. Utilizing this insight, we demonstrate that graph sparsification based on our proxy spectral gap update can perform better than or on par with a contemporary baseline (Chen et al., 2021) that takes additional node features and labels into account. This highlights the feasibility of finding winning subgraphs at initialization.
1.2 Related Work
@@
Over-squashing. Alon & Yahav (2021); Topping et al. (2022) have observed that over-squashing, where information from distant nodes are not propagated due to topological bottlenecks in the graph, hampers the performance of GNNs. A promising line of work that attempts to alleviate this issue is graph rewiring. This task aims to modify the edge structure of the graph either by adding or deleting edges. Gasteiger et al. (2019) propose to add edges according to graph diffusion kernel, such as personalized PageRank, to rely less on messages from only one-hop neighbors, thus alleviating over-squashing. Topping et al. (2022) theoretically analyze the presence of bottlenecks and find that negatively curved edges characterized by the discrete Ricci curvature (Hamilton, 1988) are the main source of information congestion. As a solution, they propose Stochastic Discrete Ricci Flow (SDRF), that surgically adds edges to support the negatively curved edges while deleting the positively curved edges. Banerjee et al. (2022) resort to measuring the spectral expansion with respect to the number of rewired edges and propose a random edge flip algorithm that transforms the given input graph into an Expander graph. Contrarily, Deac et al. (2022) show that negatively curved edges might be inevitable for building scalable GNNs without bottlenecks and advocate the use of Expander graphs for message passing.
Another line of research alleviates bottlenecks by influencing the spectral gap without taking curvature explicitly into account. Arnaiz-Rodríguez et al. (2022) introduce two new intermediate layers called CT-layer and GAP-layer, which can be interspersed between regular GNN layers. The layers perform edge re-weighting (which minimizes the gap) and introduce additional parameters. Recently, Karhadkar et al. (2023) propose FoSR, a graph rewiring algorithm that sequentially adds edges to maximize the first-order approximation of the spectral gap and also provides special labels to the newly added edges. Combined with Relational-GCN (Battaglia et al., 2018), the authors achieve performance boosts on several graph classification tasks. Giovanni et al. (2023) provide a comprehensive account of over-squashing and studies the interplay of depth, width and the topology of the graph. A recent work by Black et al. (2023) explores the idea of characterizing over-squashing through the lens of effective resistance (Chandra et al., 1996), which is inherently tied to the strength of the paths connecting two nodes in a graph. Edges are added that minimize the effective resistance of the graph and thereby mitigate over-squashing.
@@
Over-smoothing. A simple strategy to ensure GNNs can model long-range interactions is to increase the number of layers. It is a known fact that increasing network depth (He et al., 2016) often leads to better performance in the case of deep neural networks. However, naively stacking GNN layers seems to harm generalization. This problem can be seen through the lens of training dynamics. Mustafa et al. (2023) derive conservation laws for gradient flows for GATs (Veličković et al., 2018) that demonstrate the importance of healthy gradient flow in training deeper networks. This might also be approached from the perspective of over-smoothing (Li et al., 2019; Oono & Suzuki, 2020; NT & Maehara, 2019; Zhou et al., 2021; Rusch et al., 2023). In both of these cases, repeated aggregation leads to node features, in particular of nodes with different labels, becoming more similar until the nodes are hardly distinguishable. Current graph rewiring strategies, such as FoSR (Karhadkar et al., 2023), which rely on iteratively adding edges based on spectral expansion, may help mitigate over-squashing but will inevitably lead to over-smoothing. They resort to using a Relational-GCN to differentiate between newly added edges vs. existing ones to avoid strong over-smoothing. Some curvature based works (Nguyen et al., 2023; Giraldo et al., 2023) aim to optimize the degree of smoothing by graph rewiring, as they perceive over-smoothing as the result of too much information propagation, while over-squashing is caused by too little. Within this framework, they assume that edge deletions always reduce the spectral gap, while we show and exploit that some deletions can also increase it. Furthermore, we rely on a different, well established concept of over-smoothing (Keriven, 2022) that also takes node features into account and is therefore not diametrically opposed to over-squashing. In contrast to our proposal, curvature based methods (Nguyen et al., 2023; Giraldo et al., 2023) do not scale well to large graphs. For instance, Nguyen et al. (2023) propose a batch Ollivier-Ricci (BORF) curvature based rewiring approach to add and delete edges, which solves optimal transport problems and runs in cubic time.
@@
Graph sparsification and lottery tickets. Most GNNs perform recursive aggregations of neighborhood information. This operation becomes computationally expensive when the graphs are large and dense. A possible solution for this is to extract a subset of the graph which is representative of the dense graph, either in terms of their node distribution (Eden et al., 2018) or graph spectrum (Adhikari et al., 2017). Zheng et al. (2020); Li et al. (2020) formulate graph sparsification as an optimization problem by resorting to learning surrogates and ADMM respectively. With the primary aim to reduce the computational resource requirements of GNNs, a line of work that transfers the lottery ticket hypothesis (LTH) by Frankle & Carbin (2019) to GNNs (Chen et al., 2021; Hui et al., 2023), prunes the model weights in addition to the adjacency matrix. The resulting winning graph lottery ticket (GLT) can match or surpass the performance of the original dense model. While our theoretical understanding of GLTs is primarily centered around their existence (Ferbach et al., 2022; Burkholz et al., 2022; Burkholz, 2022b, a), our insights inspired by the Braess paradox add a complementary lens to our understanding how generalization can be improved, namely by reducing over-squashing and over-smoothing with graph pruning. So far, the spectral gap has only been employed to maintain a sufficient degree of connectivity of bipartite graphs that are associated with classic feed-forward neural network architectures Pal et al. (2022); Hoang et al. (2023). As we show, the spectral gap can also be employed as a pruning at initialization technique (Frankle et al., 2021) that does not take node features into account and achieves high computational resource savings. However, as we show, if pruning is gradually combined with weight sparsification during training, its generalization error can be improved, which is in line with observations for random pruning of CNNs (Gadhikar et al., 2023; Gadhikar & Burkholz, 2024).
@@ 2 Theoretical Insights Into Spectral Rewiring
To illustrate the utility of the Braess paradox, we discuss a minimal example, illustrated in Figure 1. To provide an intuition for the multiple mechanisms at work, in this setting we prove that the deletion of an edge can both alleviate over-smoothing and over-squashing. In contrast, an edge addition that addresses over-squashing still causes over-smoothing, yet less drastically than another edge addition that worsens over-squashing.
@@
Reducing over-squashing via the spectral gap. From a spectral perspective, bottlenecks, which hamper the information flow by over-squashing, can be characterized by the spectral gap of the (symmetric) normalized graph Laplacian $\mathcal{L}_{\mathcal{G}}$, where $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The Laplacian of the graph is $\mathcal{L}=D-A$, where $A$ is the adjacency matrix and $D$ the diagonal degree matrix. The symmetric normalized graph Laplacian is defined as $\mathcal{L}_{\mathcal{G}}=D^{-1/2}\mathcal{L}D^{-1/2}$. Let $\lambda_{0}<\lambda_{1}<\lambda_{2},...\lambda_{n}$} be the eigenvalues of $\mathcal{L}_{\mathcal{G}}$ arranged in ascending order and let $\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ be the first non-zero eigenvalue of the normalized graph Laplacian, which is also called the spectral gap of the graph. For a graph where distant network components are connected only by a few bridging edges, all the information has to be propagated via these edges. The information flow through edges is encoded by the Cheeger (1969) constant $h_{S}=\min_{S\subset V}\frac{|\partial S|}{\min\{Vol(S),Vol(S\backslash V)\}}$ where $\partial S=\{(u,v):u\in S,v\in\mathcal{V}\backslash S\}$ and $Vol(S)=\sum_{u\in S}d_{u}$, being $d_{u}$ the degree of the node $u$. The spectral gap is bounded by the Cheeger inequality $2h_{\mathcal{G}}\geq\lambda_{1}\geq\frac{h^{2}_{\mathcal{G}}}{2}$, which motivates it as measure of over-squashing.
Braess’ paradox. Braess (1968) found a counter-intuitive result for road networks: even if all travelers behave selfishly, the removal of a road can still improve each of their individual travel times. That is, there is a violation of monotonicity in the traffic flow with respect to the number of edges of a network. This effect is common, and not just a sporadic observation. For example, Chung & Young (2010) has shown that Braess’ paradox occurs with high probability in almost all connected Erdős-Rényi random graphs, and Chung et al. (2012) have confirmed it for a large class of expander graphs. The paradox may analogously be applied to related graph properties such as the spectral gap of the normalized Laplacian. Eldan et al. (2017) have studied how the spectral gap of a random graph changes after edge additions or deletions, proving a strictly positive occurrence of the paradox for typical instances of ER graphs. This result inspires us to develop an algorithm for rewiring a graph by specifically eliminating edges that increase this quantity, which we can expect to carry out with high confidence in real-world graphs. Their Lemma 3.2 (when reversed) states a sufficient condition that guarantees a spectral gap increase in response to a deletion of an edge.
@@
Eldan et al. (2017): Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a finite graph, with $f$ denoting the eigenvector and $\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ the eigenvalue corresponding to the spectral gap. Let $\{u,v\}\notin\mathcal{V}$ be two vertices that are not connected by an edge. Denote $-^\hat{\mathcal{G}}=(\mathcal{V},\hat{\mathcal{E}})$, the new graph obtained after adding an edge between $\{u,v\}$, i.e., ^:=$\hat{\mathcal{E}}:=\mathcal{E}\cup\{u,v\}$. Denote with $\mathcal{P}_{f}:=\langle f,\hat{f_{0}}\rangle$ the projection of $f$ onto the top eigenvector of $\hat{\mathcal{G}}$. Define $g\left(u,v,\mathcal{L}_{\mathcal{G}}\right):=$
$-\mathcal{P}^{2}_{f}\lambda_{1}(\mathcal{L}_{\mathcal{G}})-2(1-\lambda_{1}(\mathcal{L}_{\mathcal{G}}))\left(\frac{\sqrt{{d_{u}}+1}-\sqrt{d_{u}}}{\sqrt{d_{u}+1}}f_{u}^{2}\right.$
$\left.+\frac{\sqrt{{d_{v}}+1}-\sqrt{d_{v}}}{\sqrt{d_{v}+1}}f_{v}^{2}\right)+\frac{2f_{u}f_{v}}{\sqrt{d_{u}+1}\sqrt{d_{v}+1}}.$
@@
The spectral gap of $\mathcal{G}$ increases with the deletion of edge $\{0,3\}$, i.e., $\lambda_{1}(\mathcal{L}_{\mathcal{G}^{-}})>\lambda_{1}(\mathcal{L}_{\mathcal{G}})$.
It also increases with the addition of edge $\{0,5\}$ or decreases with the addition of edge $\{4,7\}$, i.e., $\lambda_{1}(\mathcal{L}_{\mathcal{G}^{+}})>\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ and $\lambda_{1}(\mathcal{L}_{\widetilde{\mathcal{G}^{+}}})<\lambda_{1}(\mathcal{L}_{\mathcal{G}})$.
The proof could just be a simple explicit calculation of the leading eigenvalues of the normalized Laplacians of the respective graphs. However, to develop an intuition and relate our contributions to this example, we leverage Eldan’s Lemma 2.1 in Appendix A.1 and apply the spectral graph proxies in our derivations starting from an explicit spectral analysis of the ring graph. While these derivations demonstrate that we can reduce over-squashing (i.e., increase the spectral gap) by edge deletions, we show next that edge deletions can also alleviate over-smoothing.
Over-smoothing. At least for GNNs with mean aggregation, increasing the spectral gap usually promotes smoothing and thus leads to higher node feature similarity. Equating a high node feature similarity with over-smoothing would thus imply a trade-off between over-smoothing and over-squashing. From this angle, some works (Giraldo et al., 2023; Nguyen et al., 2023) seek to find the right amount of smoothing by adding edges to increase or deleting edges to decrease the spectral gap. Contrarily, we argue that deleting edges can increase the gap while adding edges could decrease it, as our previous analysis demonstrates. Thus, both edge deletions and additions allow to control which node features are aggregated, while mitigating over-squashing. Such node features are central to a more nuanced concept of over-smoothing that acknowledges that increasing the similarity of nodes that share the same label, while keeping nodes with different labels distinguishable, aids the learning task.
To measure over-smoothing, we adopt the Linear GNN test bed proposed by Keriven (2022), which uses a linear ridge regression (LRR) setup with mean squared error (MSE) as the loss. We assign two classes to nodes according to their color in Figure 1, and one-dimensional features that are drawn independently from normal distributions $\mathcal{N}(1,1)$ and $\mathcal{N}(-1,1)$, respectively. Figure 2(a) compares how our exemplary graphs (see Figure 1) influence over-smoothing in this setting. While adding edges can accelerate the rate of smoothing, pruning strikingly aids in reducing over-smoothing —and still reduces over-squashing by increasing the spectral gap. Note that the real world heterophilic graph example shows a similar trend and highlights the utility of the spectral pruning algorithm ProxyDelete, which we describe in the next section, over edge additions by the strong basline FoSR. Additional real world examples can be found in Appendix D.

### v2

@@ Abstract
Message Passing Graph Neural Networks are known to suffer from two problems that are sometimes believed to be diametrically opposed: over-squashing and over-smoothing. The former results from topological bottlenecks that hamper the information flow from distant nodes and are mitigated by spectral gap maximization, primarily, by means of edge additions. However, such additions often promote over-smoothing that renders nodes of different classes less distinguishable. Inspired by the Braess phenomenon, we argue that deleting edges can address over-squashing and over-smoothing simultaneously. This insight explains how edge deletions can improve generalization, thus connecting spectral gap optimization to a seemingly disconnected objective of reducing computational resources by pruning graphs for lottery tickets. To this end, we propose a computationally effective spectral gap optimization framework to add or delete edges and demonstrate its effectiveness on the long range graph benchmark and on larger heterophilous datasets.
@@ 1 Introduction
Graphs are ubiquitous data structures that can model data from diverse fields ranging from chemistry (Reiser et al., 2022), biology (Bongini et al., 2023) to even high-energy physics (Shlomi et al., 2021). This has led to the development of deep learning techniques for graphs, commonly referred to as Graph Neural Networks (GNNs). The most popular GNNs follow the message-passing paradigm (Gori et al., 2005; Scarselli et al., 2009; Gilmer et al., 2017; Bronstein et al., 2021), where arbitrary differentiable functions, parameterized by neural networks, are used to diffuse information on the graph, consequently learning a graph-level representation. This representation can then be used for various downstream tasks like node classification, link prediction, and graph classification. Different types of GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Xu et al., 2019; Bodnar et al., 2021a, b; Bevilacqua et al., 2022), all tackling a variety of problems in various domains have been proposed with varied degree of success. Despite their widespread use, GNNs have a number of inherent problems. These include limited expressivity, (Leman, 1968; Morris et al., 2019), over-smoothing (Li et al., 2019; NT & Maehara, 2019; Oono & Suzuki, 2020; Zhou et al., 2021), and over-squashing (Alon & Yahav, 2021; Topping et al., 2022).
@@
The phenomenon of over-squashing, first studied heuristically by Alon & Yahav (2021) and later theoretically formalized by Topping et al. (2022), is caused by the presence of structural bottlenecks in the graph. These bottlenecks can be attributed to the first non-zero eigenvalue of the normalized graph Laplacian, also known as the spectral gap. The smaller the gap, the more susceptible a graph is to over-squashing. Recent work has explored rewiring the input graph to address these bottlenecks (Topping et al., 2022; Arnaiz-Rodríguez et al., 2022; Giraldo et al., 2023; Nguyen et al., 2023; Karhadkar et al., 2023), but suggest there has to be a trade-off between over-squashing and over-smoothing (Keriven, 2022). Instead, we propose to leverage the Braess paradox (Braess, 1968; Eldan et al., 2017) that posits certain edge deletions can maximize the spectral gap. We propose to approximate the spectral change in a computationally efficient manner by leveraging Matrix Perturbation Theory (Stewart & Sun, 1990). Our proposed framework allows us to jointly address the problem of over-squashing, by increasing the spectral gap, and over-smoothing, by slowing down the rate of smoothing. We find that our method is especially effective in heterophilic graph settings, where we delete edges between nodes of different labels, thus preventing unnecessary aggregation. We empirically show that our proposed method outperforms other graph rewiring methods on node classification and graph classification tasks. We also show that spectral gap based edge deletions can help identify graph lottery tickets (GLTs) (Frankle & Carbin, 2019), that is, sparse sub-networks that can match the performance of dense networks.
@@ 1.1 Contributions
  1. 1.
    Inspired by the Braess phenomenon, we prove that, contrary to common assumptions, over-smoothing and over-squashing are not necessarily diametrically opposed. By deriving a minimal example, we show that both can be mitigated by spectral based edge deletions.
  2. 2.
    Leveraging matrix perturbation theory, we propose a Greedy graph pruning algorithm (ProxyDelete) that maximizes the spectral gap in a computationally efficient way. Similarly, our algorithm can also be utilized to add edges in a joint framework. We compare this approach with a novel graph rewiring scheme based on Eldan’s criterion (Eldan et al., 2017) that provides guarantees for edge deletions and a stopping criterion for pruning, but is computationally less efficient.
  3. 3.
@@
    Our results connect literature on three seemingly disconnected topics: over-smoothing, over-squashing, and graph lottery tickets, which explain observed improvements in generalization performance by graph pruning. Utilizing this insight, we demonstrate that graph sparsification based on our proxy spectral gap update can perform better than or on par with a contemporary baseline (Chen et al., 2021) that takes additional node features and labels into account. This highlights the feasibility of finding winning subgraphs at initialization.
2 Related work
@@
Over-squashing. Alon & Yahav (2021); Topping et al. (2022) have observed that over-squashing, where information from distant nodes are not propagated due to topological bottlenecks in the graph, hampers the performance of GNNs. A promising line of work that attempts to alleviate this issue is graph rewiring. This task aims to modify the edge structure of the graph either by adding or deleting edges. Gasteiger et al. (2019) propose to add edges according to graph diffusion kernel, such as personalized PageRank, to rely less on messages from only one-hop neighbors, thus alleviating over-squashing. Topping et al. (2022) propose Stochastic Discrete Ricci Flow (SDRF) to rewire the graph based on curvature. Banerjee et al. (2022) resort to measuring the spectral expansion with respect to the number of rewired edges and propose a random edge flip algorithm that transforms the given input graph into an Expander graph. Contrarily, Deac et al. (2022) show that negatively curved edges might be inevitable for building scalable GNNs without bottlenecks and advocate the use of Expander graphs for message passing. Arnaiz-Rodríguez et al. (2022) introduces two new intermediate layers called CT-layer and GAP-layer, which can be interspersed between GNN layers. The layers perform edge re-weighting (which minimizes the gap) and introduce additional parameters. Karhadkar et al. (2023) propose FoSR, a graph rewiring algorithm that sequentially adds edges to maximize the first-order approximation of the spectral gap. A recent work by Black et al. (2023) explores the idea of characterizing over-squashing through the lens of effective resistance (Chandra et al., 1996). Giovanni et al. (2023) provide a comprehensive account of over-squashing and studies the interplay of depth, width and the topology of the graph.
@@
Over-smoothing. It is a known fact that increasing network depth (He et al., 2016) often leads to better performance in the case of deep neural networks. However, naively stacking GNN layers often seems to harm generalization. And one of the reasons is over-smoothing (Li et al., 2019; Oono & Suzuki, 2020; NT & Maehara, 2019; Zhou et al., 2021; Rusch et al., 2023a), where repeated aggregation leads to node features, in particular nodes with different labels, becoming indistinguishable. Current graph rewiring strategies, such as FoSR (Karhadkar et al., 2023), which rely on iteratively adding edges based on spectral expansion, may help mitigate over-squashing but also increase the smoothing induced by message passing. Curvature based methods such as Nguyen et al. (2023); Giraldo et al. (2023) aim to optimize the degree of smoothing by graph rewiring, as they assume that over-smoothing is the result of too much information propagation, while over-squashing is caused by too little. Within this framework, they assume that edge deletions always reduce the spectral gap. In contrast, we show and exploit that some deletions can also increase it. Furthermore, we rely on a different, well established concept of over-smoothing (Keriven, 2022) that also takes node features into account and is therefore not diametrically opposed to over-squashing. As we show, over-smoothing and over-squashing can be mitigated jointly. Moreover, we propose a computationally efficient approach to achieve this with spectral rewiring. In contrast to our proposal, curvature based methods (Nguyen et al., 2023; Giraldo et al., 2023) do not scale well to large graphs. For instance, Nguyen et al. (2023) propose a batch Ollivier-Ricci (BORF) curvature based rewiring approach to add and delete edges, which solves optimal transport problems and runs in cubic time.
@@
Graph sparsification and lottery tickets. Most GNNs perform recursive aggregations of neighborhood information. This operation becomes computationally expensive when the graphs are large and dense. A possible solution for this is to extract a subset of the graph which is representative of the dense graph, either in terms of their node distribution (Eden et al., 2018) or graph spectrum (Adhikari et al., 2017). Zheng et al. (2020); Li et al. (2020) formulate graph sparsification as an optimization problem by resorting to learning surrogates and ADMM respectively. With the primary aim to reduce the computational resource requirements of GNNs, a line of work that transfers the lottery ticket hypothesis (LTH) by Frankle & Carbin (2019) to GNNs (Chen et al., 2021; Hui et al., 2023), prunes the model weights in addition to the adjacency matrix. The resulting winning graph lottery ticket (GLT) can match or surpass the performance of the original dense model. While our theoretical understanding of GLTs is primarily centered around their existence (Ferbach et al., 2022; Burkholz et al., 2022; Burkholz, 2022b, a), our insights inspired by the Braess paradox add a complementary lens to our understanding of how generalization can be improved, namely by reducing over-squashing and over-smoothing with graph pruning. So far, the spectral gap has only been employed to maintain a sufficient degree of connectivity of bipartite graphs that are associated with classic feed-forward neural network architectures (Pal et al., 2022; Hoang et al., 2023). We highlight that the spectral gap can also be employed as a pruning at initialization technique (Frankle et al., 2021) that does not take node features into account and can achieve computational resource savings while reducing the generalization error, which is in line with observations for random pruning of CNNs (Gadhikar et al., 2023; Gadhikar & Burkholz, 2024).
@@ 3 Theoretical insights into spectral rewiring
To prove our claim that over-smoothing and over-squashing can both be alleviated jointly, we provide a minimal example as illustrated in Figure 1. Utilizing the Braess paradox, we achieve this by the deletion of an edge. In contrast, an edge addition that addresses over-squashing still causes over-smoothing, yet less drastically than another edge addition that worsens over-squashing.
Reducing over-squashing via the spectral gap. From a spectral perspective, bottlenecks, which hamper the information flow by over-squashing, can be characterized by the spectral gap of the (symmetric) normalized graph Laplacian $\mathcal{L}_{\mathcal{G}}$, where $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The Laplacian of the graph is $\mathcal{L}=D-A$, where $A$ is the adjacency matrix and $D$ the diagonal degree matrix. The symmetric normalized graph Laplacian is defined as $\mathcal{L}_{\mathcal{G}}=D^{-1/2}\mathcal{L}D^{-1/2}$. Let $\lambda_{0}<\lambda_{1}<\lambda_{2},...\lambda_{n}$} be the eigenvalues of $\mathcal{L}_{\mathcal{G}}$ arranged in ascending order and let $\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ be the first non-zero eigenvalue of the normalized graph Laplacian, which is also called the spectral gap of the graph. For a graph where distant network components are connected only by a few bridging edges, all the information has to be propagated via these edges. The information flow through edges is encoded by the Cheeger (1971) constant $h_{S}=\min_{S\subset V}\frac{|\partial S|}{\min\{Vol(S),Vol(S\backslash V)\}}$ where $\partial S=\{(u,v):u\in S,v\in\mathcal{V}\backslash S\}$ and $Vol(S)=\sum_{u\in S}d_{u}$, being $d_{u}$ the degree of the node $u$. The spectral gap is bounded by the Cheeger inequality $2h_{\mathcal{G}}\geq\lambda_{1}\geq\frac{h^{2}_{\mathcal{G}}}{2}$, which motivates it as a measure of over-squashing.
Braess’ paradox. Braess (1968) found a counter-intuitive result for road networks: even if all travelers behave selfishly, the removal of a road can still improve each of their individual travel times. That is, there is a violation of monotonicity in the traffic flow with respect to the number of edges of a network. For instance, Chung & Young (2010) has shown that Braess’ paradox occurs with high probability in Erdős-Rényi random graphs, and Chung et al. (2012) have confirmed it for a large class of Expander graphs. The paradox can be analogously applied to related graph properties such as the spectral gap of the normalized Laplacian. Eldan et al. (2017) have studied how the spectral gap of a random graph changes after edge additions or deletions, proving a strictly positive occurrence of the paradox for typical instances of ER graphs. This result inspires us to develop an algorithm for rewiring a graph by specifically eliminating edges that increase this quantity, which we can expect to carry out with high confidence in real-world graphs. Their Lemma 3.2 (when reversed) states a sufficient condition that guarantees a spectral gap increase in response to a deletion of an edge.
@@
Eldan et al. (2017): Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a finite graph, with $f$ denoting the eigenvector and $\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ the eigenvalue corresponding to the spectral gap. Let $\{u,v\}\notin\mathcal{V}$ be two vertices that are not connected by an edge. Denote $\hat{\mathcal{G}}=(\mathcal{V},\hat{\mathcal{E}})$, the new graph obtained after adding an edge between $\{u,v\}$, i.e., ^:=$\hat{\mathcal{E}}:=\mathcal{E}\cup\{u,v\}$. Denote with $\mathcal{P}_{f}:=\langle f,\hat{f_{0}}\rangle$ the projection of $f$ onto the top eigenvector of $\hat{\mathcal{G}}$. Define $g\left(u,v,\mathcal{L}_{\mathcal{G}}\right):=$
$-\mathcal{P}^{2}_{f}\lambda_{1}(\mathcal{L}_{\mathcal{G}})-2(1-\lambda_{1}(\mathcal{L}_{\mathcal{G}}))\left(\frac{\sqrt{{d_{u}}+1}-\sqrt{d_{u}}}{\sqrt{d_{u}+1}}f_{u}^{2}\right.\left.+\frac{\sqrt{{d_{v}}+1}-\sqrt{d_{v}}}{\sqrt{d_{v}+1}}f_{v}^{2}\right)+\frac{2f_{u}f_{v}}{\sqrt{d_{u}+1}\sqrt{d_{v}+1}}.$
@@
The spectral gap of $\mathcal{G}$ increases with the deletion of edge $\{0,3\}$, i.e., $\lambda_{1}(\mathcal{L}_{\mathcal{G}^{-}})>\lambda_{1}(\mathcal{L}_{\mathcal{G}})$. It also increases with the addition of edge $\{0,5\}$ or decreases with the addition of edge $\{4,7\}$, i.e., $\lambda_{1}(\mathcal{L}_{\mathcal{G}^{+}})>\lambda_{1}(\mathcal{L}_{\mathcal{G}})$ and $\lambda_{1}(\mathcal{L}_{\widetilde{\mathcal{G}^{+}}})<\lambda_{1}(\mathcal{L}_{\mathcal{G}})$.
We leverage Eldan’s Lemma 3.1 in Appendix A.1 and apply the spectral graph proxies in our derivations starting from an explicit spectral analysis of the ring graph. While these derivations demonstrate that we can reduce over-squashing (i.e., increase the spectral gap) by edge deletions, we show next that edge deletions can also alleviate over-smoothing.
Slowing detrimental over-smoothing. For GNNs with mean aggregation, increasing the spectral gap usually promotes smoothing and thus leads to higher node feature similarity. Equating a high node feature similarity with over-smoothing would thus imply a trade-off between over-smoothing and over-squashing. Methods by Giraldo et al. (2023); Nguyen et al. (2023) seek to find the right amount of smoothing by adding edges to increase the gap and deleting edges to decrease it. Contrarily, we argue that deleting edges can also increase the gap while adding edges could decrease it, as our previous analysis demonstrates. Thus, both edge deletions and additions allow to control which node features are aggregated, while mitigating over-squashing. Such node features are central to a more nuanced concept of over-smoothing that acknowledges that increasing the similarity of nodes that share the same label, while keeping nodes with different labels distinguishable, aids the learning task.
To measure over-smoothing, we adopt the Linear GNN test bed proposed by Keriven (2022), which uses a linear ridge regression (LRR) setup with mean squared error (MSE) as the loss. We assign two classes to nodes according to their color in Figure 1, and one-dimensional features that are drawn independently from normal distributions $\mathcal{N}(1,1)$ and $\mathcal{N}(-1,1)$, respectively. Figure 2(a) compares how our exemplary graphs (see Figure 1) influence over-smoothing in this setting. While adding edges can accelerate the rate of smoothing, pruning strikingly aids in reducing over-smoothing —and still reduces over-squashing by increasing the spectral gap. Note that the real world heterophilic graph example shows a similar trend and highlights the utility of the spectral pruning algorithm ProxyDelete, which we describe in the next section, over edge additions by the strong baseline FoSR. Additional real world examples along with cosine distance between nodes of different labels before and after spectral pruning and plots for Dirichlet energy can be found in Appendix D.